a 50 kg crate is pushed by a 1n applied force and pulled by a rope with make 40 with the horizontal if the Lensim is 5n
A. find the normal force
B. Find friction force

Given.
Fa =1n
T=5n,40n of e
M=50 kg
1
Expert's answer
2019-11-25T10:08:27-0500

A. Normal force:


"N=mg-T\sin{40}"

"N=(50)(9.8)-5\sin{40}=487\ N"

B. Friction force:


"F_f=T\cos{40}-F_a"

"F_f=5\cos{40}-1=2.83\ N"